OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? A triangle with vertices at (1, 2), (–1, –6) and (7, –2) is translated five units to the left and 7 units upward. The coordinates of the translated triangle are: a. (6, 9), (4, 1) and (12, 5) b. (–4, –5), (–6, –13) and (2, –9) c. (6, –5), (4, –13) and (12, –9) d. (–4, 9), (–6, 1) and (2, 5)

OpenStudy (igreen):

5 units left will be - 5 for the x-value 7 units upward will be + 7 for the y-value Just subtract 5 from x and add 7 to y

OpenStudy (igreen):

So: 1 - 5 = ? 2 + 7 = ?

OpenStudy (igreen):

@ninaesb

OpenStudy (anonymous):

@iGreen 1-5=-4 2+7=9

OpenStudy (igreen):

Right, so the first coordinate is (-4, 9) and only one Option has that..which one is it?

OpenStudy (igreen):

@ninaesb

OpenStudy (anonymous):

D :) i got it . Thanks @iGreen

OpenStudy (igreen):

Yep!